OpenStudy (anonymous):

I have a question that says a rocket is launced from top of a a80 ft cliff with a initial celocity of 140 ft per second. the height of h of the orcket after t seconds is given by the euqation-16t^2+140t+80. how long after the rocket is launced will it be 20 ft from the ground. I have tried several problems like this and I still don't get it. can you help?

OpenStudy (anonymous):

set equation = 20 and solve for t

OpenStudy (anonymous):

so how do you do this?

OpenStudy (anonymous):

\[-16t^2+140t+80=20\] i think there was a typo in the problem because it should be \[-16t^2\] and not \[16t^2\]

OpenStudy (anonymous):

yes it should be -16t^2

OpenStudy (anonymous):

can you walk me through the step so that I understand it?

OpenStudy (anonymous):

\[-16t^2+140t+80=20\] \[4t^2-35t-15=0\] then probably quadratic formula

OpenStudy (anonymous):

They are using the sq rt and I do not understand this?

OpenStudy (anonymous):

What do you mean by "using sqrt"?

OpenStudy (anonymous):

r I see you do not understand it either so I will try somewhere else

OpenStudy (anonymous):

:)